scenario

You are a marketing consultant for a specialty coffee bistro and bakery. The business has been operating for about a year and is still not earning a profit. Your role is to determine the breakeven point and make recommendations on changes to one or more of the 4 P's: Product, Price, Promotion, and Place.

Breakeven point

The breakeven point (BEP) is the point at which total cost and total revenue are equal. A breakeven analysis calculates the BEP as:

Breakeven Point = Total Fixed Costs / (Unit Price - Unit Variable Cost) or BEP = F / (P-V)

In our example, the BEP = Total Fixed Costs / (Average Order Price - Variable Cost per Order). Using this formula, what is the breakeven point? In other words, how many orders would need to be sold before the business starts making a profit given the current price and cost structure?

Here's the information you need to do the bistro's BEP analysis:

Current Orders per Month

750

Average Order Price

$11.00

Variable Cost per Order

$5.00

Rent per Month

$2,000.00

Utilities per Month

$400.00

Labor per Month

$3,000.00

Calculate the BEP using the numbers and formulas above. Hint: total fixed costs in this case are the sum of rent, utilities and labor.

Recommendations

In order to be profitable, the business will need to either increase revenue or decrease expenses. Make three recommendations on changes to the Product, Price, Promotion and/or Place that would be likely to increase revenue, decrease expenses, or both. Be creative and have fun with your ideas!

Here are some examples of questions you could ask to help with your analysis:

  • What changes could be made to the product to reduce variable costs?
  • What changes could be made to prices to increase revenue, and how would you justify those price changes?
  • Is increasing price the only viable pricing strategy, or could lower prices lead to increased orders beyond the bistro's BEP?
  • What changes could be made to promotions (e.g., advertising, discounts, sales, loyalty programs) to increase the number of orders per month, remembering that promotions are expenses?

After you calculate BEP, make three recommendations and be sure to explain why you believe your recommended changes could achieve the business objectives of reaching profitability by either increasing revenue or decreasing expenses, and recalculate the BEP based upon your recommendations to ensure that your recommendations result in profitability.

Assignment

Create a Thread presentation with comment that includes your BEP calculations and three recommendations on changes to the 4 P's:

  • Slide 1 - Introduction: include your name, instructor's name, course, and assignment due date
  • Slide 2 - Breakeven analysis
  • Slide 3 - Three recommendations for changes to the 4 P's

please help!!!
Answer #1

Calculation of Breakeven Point of the specialty coffee bistro and bakery company:

Given,

Average order Price = $ 11

Variable cost per order = $ 5

Rent per month =$ 2000

Utilities per month = $ 400

Labor per month = $ 3000

Total Fixed Cost = Rent + Utilities + Labor Cost =$ (2000+400+3000) = $ 5400

Breakeven Point = Total Fixed Cost/( Average order Price - Variable cost per order) = [5400 / (11-5)] units = 900 units

Given, Current order per month = 750 units

So, the company is currently producing (900 -750) or 150 units less to reach the breakeven point.

Recommendation 1:

The company can increase the Order Price per unit to reach the breakeven point, considering the current order units, variable price per unit and fixed cost constant.

Let, the new order price = a

So, as per Breakeven formula, [5400 / (a-5)] units = 750 units

Therefore, a =12.2

If the company’s average order price increased to $ 12.2 from current $ 11; it will reach the breakeven point.

Recommendation 2:

The company can decrease the variable price per unit to reach the breakeven point, considering the current order units, order price per unit and fixed cost constant.

Let, the new order price per unit = b

So, as per Breakeven formula, [5400 / (11-b)] units = 750 units

Therefore, b =3.8

If the company’s variable price per unit decreased to $ 3.8 from current $ 5; it will reach the breakeven point.

Recommendation 3:

The company may go for Promotional activities like advertising, discounts, sales, loyalty programs etc. Though in that case, the total fixed cost will increase; but they can anticipate how many new orders can be drawn from the promotional activity. By using the same formula, they can also predict those promotional activities are viable or not and can choose the correct one.